# Are penalty shootouts better than a coin toss? Evidence from international club football in Europe

### v4

@@ Abstract
Penalty shootouts play a crucial role in the knockout stage of major football tournaments. Their importance has been substantially increased from the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule. Our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, we do not detect any (positive) relationship between relative team strength and shootout success using differences in Elo ratings. Consequently, penalty shootouts seem to be close to a coin toss in top European club football.
@@ 1 Introduction
As we demonstrate in Section 2, penalty shootouts are the subject of serious academic interest in the last one and a half decades. The present paper contributes to this discussion by investigating the following research question: Does the shooting order, the field of the match, psychological momentum, or the strength of the teams affect the outcome of a penalty shootout? All of these issues have partially been investigated in the existing literature as presented in Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, the psychological momentum enjoyed by the comeback team is identified in two ways, and one of them—which team scored the last goal?—is novel. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose.
@@
As we show, the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025 appear to approximate a coin toss. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), which cannot be observed in UEFA club competitions using Football Club Elo Ratings. A possible reason is the smaller strength difference in UEFA club competitions compared to national cups, which may limit the detectable effect of strength. Naturally, our finding cannot be interpreted as “ability does not matter”: the analysis is conditional on matches reaching a penalty shootout, and stronger teams have recently been demonstrated to have a higher probability of qualifying in UEFA club competitions (Csató,, 2024).
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15% (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since the outcome of a penalty shootout appears to be close to a coin toss in UEFA club competitions with respect to these observables, such a measure might favour weaker teams and teams playing away that, contrary to extra time (see Bahamonde-Birke and Bahamonde-Birke, (2023) for the latter effect), are not disadvantaged in the shootout, level the playing field, and decrease the dominance of the top clubs.
@@
The paper is structured as follows. An overview of related studies is given in Section 2. The data and the methodology are described in Sections 3 and 4, respectively. Section 5 presents and discusses the results, while Section 6 concludes.
@@ 2 Related literature
The issue of first-mover advantage in football penalty shootouts has received serious attention since the pioneering work of Apesteguia and Palacios-Huerta, (2010). Some studies find evidence for a significant advantage enjoyed by the team kicking the first penalty in each round (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020). These results have inspired formal modelling of the first-mover advantage, as well as several proposals for alternative mechanisms to mitigate or even eliminate this source of unfairness (Palacios-Huerta,, 2012; Echenique,, 2017; Brams and Ismail,, 2018; Vandebroek et al.,, 2018; Del Giudice,, 2019; Anbarcı et al.,, 2021; Csató, 2021a, ; Csató, 2021b, ; Csató and Petróczy,, 2022; Brams et al.,, 2024). The alternating (or ABBA) order was even tried in various tournaments between 2017 and 2018 (Palacios-Huerta,, 2020; Csató, 2021a, ), albeit the IFAB (International Football Association Board), the rule-making body of football, stopped the experiments in 2018 (FIFA,, 2018). It is worth noting that the first-mover advantage seems to be non-existent in tennis tiebreaks, which use the ABBA sequence (Cohen-Zada et al.,, 2018; Da Silva et al.,, 2018).
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023). According to Kassis et al., (2021), teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. These conflicting results can partially be attributed to limited sample sizes. According to the simulation exercise of Vandebroek et al., (2018), a sample of 540 shootouts with a fixed underlying mathematical model produces considerable variation: the winning proportion of a team that enjoys a significant advantage in the model becomes significant in only 56.5% of the 10 thousand simulations at 5% level of significance.
@@
Regarding the existence of a penalty-specific home advantage, Apesteguia and Palacios-Huerta, (2010) show no difference in the chances of home and away teams for 129 shootouts between 1976 and 2003. Kocher et al., (2012) reinforce this based on 540 shootouts between 1970 and 2003. Analogously, playing at home does not influence the probability of winning in 252 shootouts from French cup competitions (Arrondel et al.,, 2019). The results of Wunderlich et al., (2020) suggest no home advantage in 1067 penalty shootouts that took place between 2004 and 2018. The logistic regression of Bahamonde-Birke and Bahamonde-Birke, (2023) based on a sample of 471 shootouts demonstrates the absence of such an effect, too. Even though van Ours and van Tuijl, (2024) state that a penalty shootout does not seem to be a lottery in the Dutch domestic cup, the 20 home wins out of the 33 shootouts (61%) are not significantly different from 50%.
We know only one study on the impact of psychological momentum, although the performance of men is significantly affected by psychological momentum in sports (Cohen-Zada et al.,, 2017). Krumer, (2021) analyses 214 penalty shootouts from two-legged matches in European cups (European Champion Clubs’ Cup/UEFA Champions League, European Cup Winners’ Cup/UEFA Cup Winners’ Cup, UEFA Cup/UEFA Europa League) between 1970 and 2018. Every additional goal scored by the away team in regular time in the second leg increases its probability of winning the penalty shootout by 12.5 percentage points. However, if the home team wins in regular time, it has only a 50% chance to qualify after a shootout.
@@
Stronger football teams have a significantly higher chance of winning a penalty shootout based on 1067 shootouts from the domestic cup competitions of ten European associations (the top five, as well as Belgium, Portugal, Russia, Turkey, Ukraine), the Brazilian cup, the UEFA Champions League, and the UEFA Europa League (Wunderlich et al.,, 2020). However, the effect of team strength on success remains rather small; the winning probability does not exceed 60% even against an extremely weak team. The novelty of Wunderlich et al., (2020) resides in the measure of strength difference, determined by winning probabilities based on pre-match betting odds, adjusted for home advantage. The recent analysis of all available penalty shootouts until the 2023/24 season reinforces that the favourites identified by pre-match odds are more likely to win (Pipke,, 2025, Table A.5). But the favourite team might not be the stronger team due to home advantage (Wunderlich et al.,, 2020).
@@
To summarise, all studies considering the strength of teams find at least some bias towards the stronger team in winning penalty shootouts. In other words, penalty shootouts are not equivalent to a pure coin toss; ability is an important success factor even for a mechanical task carried out in a high-pressure environment.
@@ 3 Data
Champions League Europa League Conference League Total Qualification 41 (19) 103 (31) 63 (63) 207 (113) Knockout stage 14 (5) 27 (9) 9 (9) 50 (23) Final 6 (0) 5 (3) 0 (0) 11 (3) Total 61 (24) 135 (43) 72 (72) 268 (139) UEFA Conference League was called UEFA Europa Conference League until the 2023/24 season. The row Knockout stage counts only two-legged matches, but not the final, which is played on a neutral field. Numbers in parenthesis show the matches played from the 2020/21 season to 2025.
@@
Furthermore, the betting odds published by bookmakers are usually related to the outcome of one particular match, but our database mostly contains the second legs of knockout ties. Therefore, the favourite team identified by betting odds might not be the stronger team because of home advantage. Even though Wunderlich et al., (2020, Section 2.3) suggests a method to adjust this bias, which allows the classification of a slight favourite as the slightly weaker team, such a procedure is not necessarily reliable in two-legged matches. The reason is the incentives of the teams: one team could be satisfied by a small loss if it still implies its qualification. For instance, if the away team won the first match by two goals, it does not have powerful incentives to attack in the second match even if it has scored one goal less than the other team. The betting odds of a home/away win certainly reflect the effect of these incentives, which have substantially changed from the 2021/22 season due to the abolition of the away goals rule.
Compared to Elo ratings, betting odds can indeed reflect better the current form of the teams by taking, for example, injuries, into account. However, it is far from clear that these factors are relevant to penalty shootouts, where success is based on individual skills, not on collaboration within a team. The use of Elo ratings, given by an “objective” mathematical formula, may also be more persuasive for decision-makers than betting odds driven by market sentiment. Finally, Football Club Elo Ratings have recently been demonstrated to outperform UEFA club coefficients, the official measure of team strength in these competitions, with respect to predictive accuracy (Csató,, 2024). Based on these arguments, we follow the approach of previous papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), and do not consider alternative measures of team ability.
@@ 4 Methodology
The connection between the relative strength of teams and the probability of winning the penalty shootout is assessed by logistic regressions, too. The binary dependent variable is the home win/loss. Therefore, matches played on a neutral venue are removed due to the small sample size, similar to (Wunderlich et al.,, 2020). Denote the Elo rating of the home and the away team by $E_{h}$ and $E_{a}$, respectively. The independent variable $\Delta_{ha}$ is either the difference of Elo ratings ($\Delta_{ha}^{(1)}=E_{h}-E_{a}$), or the winning probability of the home team implied by the Elo ratings according to the Elo equation of Football Club Elo Ratings (http://clubelo.com/system):
@@
$P_{h}=\frac{1}{1+e^{\beta_{0}+\beta_{1}\Delta+\beta_{2}f_{h}+\beta_{3}c_{h}}},$
@@
5 Results and discussion
Section 5.1 investigates potential success factors in penalty shootouts by binomial tests, that is, whether teams in a given category win significantly more/less often than 50% compared to teams in the other category. Section 5.2 considers more dimensions simultaneously via logistic regressions. In both cases, sensitivity analyses are also conducted. Finally, we provide some explanations for the results in Section 5.3.
@@ 5.1 Partial analysis
Figure 4 focuses on the winning probability of the stronger team based on Football Club Elo Ratings. Five definitions of the favourite are used by gradually restricting the sample according to the minimal difference in the Elo ratings of the two teams. However, even teams that have at least 100 points more than their opponents fail to win 50% of their penalty shootouts. These calculations can be reproduced on the basis of Table A.2 in Appendix A.2, which also reports the associated $p$-values. The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025) demonstrate a significantly higher winning probability for a “sufficiently” stronger team.
As we have seen in Section 3, the number of penalty shootouts has substantially increased from the 2020/21 season. Thus, all binomial tests have also been conducted for the first 20 and the last six seasons, respectively. According to Tables A.1A.2 in the Appendix, the conclusions remain unchanged if we focus on the beginning or the end of the dataset. The only significant value at the level of 5% (but not at the level of 1%) appears for home advantage based on the extended dataset: home teams are less likely to win a penalty shootout than their opponents between 2020 and 2025 if the matches played behind closed doors are taken into account (Table A.1). In these 20 one-legged matches, only seven were won by the home team, which is significantly below 50% at a significance level of 10%. This can be explained by choking under pressure (Harb-Wu and Krumer,, 2019), which was probably much stronger in these matches than usual, despite the absence of spectators, as, in contrast to the standard rule, the randomly drawn home team should not have played away. The marginally significant effect disappears when only proper two-legged matches are considered under the restricted definition of home venue (Table A.1). Tables A.3 and A.4 reinforce the insignificant impact of all factors considered by distinguishing the penalty shootouts played in the qualification and knockout phases.

### v5

@@ Abstract
Penalty shootouts play a crucial role in the knockout stage of major football tournaments. Their importance has substantially increased from the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule. Our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, no evidence is found for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, we do not detect any relationship between shootout success and relative team strength, quantified by differences in Elo ratings and the implied winning probability. Thus, the hypothesis that penalty shootouts are close to a coin toss in international competitions for European football clubs cannot be rejected.
@@ 1 Introduction
As we demonstrate in Section 2, penalty shootouts are the subject of serious academic interest in the last one and a half decades. The present paper contributes to this discussion by investigating the following research question: Does the shooting order, the field of the match, psychological momentum, or the strength of the teams affect the outcome of a penalty shootout? All of these issues have partially been investigated in the existing literature as presented in Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, the psychological momentum enjoyed by the comeback team is identified in two ways, and one of them—which team scored the last goal?—is novel, even though it is also included in a paper written simultaneously with our study (van Hemert and van der Kamp,, 2026). Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), betting odds (Wunderlich et al.,, 2020; Pipke,, 2025), or (relative) market values (van Hemert and van der Kamp,, 2026) for this purpose.
@@
In our sample the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025, none of the variables above has a significant effect on winning. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025; van Hemert and van der Kamp,, 2026), which cannot be observed in UEFA club competitions using Football Club Elo Ratings. A possible reason is the smaller strength difference in UEFA club competitions compared to national cups, which may limit the detectable effect of strength. Naturally, our finding cannot be interpreted as “ability does not matter”: the analysis is conditional on matches reaching a penalty shootout, and stronger teams have recently been demonstrated to have a higher probability of qualifying in UEFA club competitions (Csató,, 2024).
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15%, in an ex ante perfectly balanced tie (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since the outcome of a penalty shootout appears to be close to a coin toss in UEFA club competitions with respect to these observables, such a measure might favour weaker teams and teams playing away that, contrary to extra time (see Bahamonde-Birke and Bahamonde-Birke, (2023) for the latter effect), are not disadvantaged in the shootout, level the playing field, and decrease the dominance of the top clubs.
@@
The paper is structured as follows. An overview of related studies is given in Section 2. The data and the methodology are described in Sections 3 and 4, respectively. Section 5 presents our results, while Section 6 offers their discussion and some concluding remarks.
@@ 2 Related literature
The issue of first-mover advantage in football penalty shootouts has received serious attention since the pioneering work of Apesteguia and Palacios-Huerta, (2010). Some studies find evidence for a significant advantage enjoyed by the team kicking the first penalty in each round (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020). These results have inspired formal modelling of the first-mover advantage, as well as several proposals for alternative mechanisms to mitigate or even eliminate this source of unfairness (Palacios-Huerta,, 2012; Echenique,, 2017; Brams and Ismail,, 2018; Vandebroek et al.,, 2018; Del Giudice,, 2019; Anbarcı et al.,, 2021; Csató, 2021a, ; Csató, 2021b, ; Csató and Petróczy,, 2022; Brams et al.,, 2024; Huang et al.,, 2026). The alternating (or ABBA) order was even tried in various tournaments between 2017 and 2018 (Palacios-Huerta,, 2020; Csató, 2021a, ), albeit the IFAB (International Football Association Board), the rule-making body of football, stopped the experiments in 2018 (FIFA,, 2018). It is worth noting that the first-mover advantage seems to be non-existent in tennis tiebreaks, which use the ABBA sequence (Cohen-Zada et al.,, 2018; Da Silva et al.,, 2018).
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023; van Hemert and van der Kamp,, 2026). According to Kassis et al., (2021), teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. These conflicting results can partially be attributed to limited sample sizes. According to the simulation exercise of Vandebroek et al., (2018), a sample of 540 shootouts with a fixed underlying mathematical model produces considerable variation: the winning proportion of a team that enjoys a significant advantage in the model becomes significant in only 56.5% of the 10 thousand simulations at 5% level of significance.
@@
Regarding the existence of a penalty-specific home advantage, Apesteguia and Palacios-Huerta, (2010) show no difference in the chances of home and away teams for 129 shootouts between 1976 and 2003. Kocher et al., (2012) reinforce this based on 540 shootouts between 1970 and 2003. Analogously, playing at home does not influence the probability of winning in 252 shootouts from French cup competitions (Arrondel et al.,, 2019). The results of Wunderlich et al., (2020) suggest no home advantage in 1067 penalty shootouts that took place between 2004 and 2018. The logistic regressions of Bahamonde-Birke and Bahamonde-Birke, (2023) and van Hemert and van der Kamp, (2026), based on 471 and 409 shootouts, respectively, demonstrate the absence of such an effect, too. Even though van Ours and van Tuijl, (2024) state that a penalty shootout does not seem to be a lottery in the Dutch domestic cup, the 20 home wins out of the 33 shootouts (61%) are not significantly different from 50%.
We know only two studies on the impact of psychological momentum, although the performance of men is significantly affected by psychological momentum in sports (Cohen-Zada et al.,, 2017). Krumer, (2021) analyses 214 penalty shootouts from two-legged matches in European cups (European Champion Clubs’ Cup/UEFA Champions League, European Cup Winners’ Cup/UEFA Cup Winners’ Cup, UEFA Cup/UEFA Europa League) between 1970 and 2018. Every additional goal scored by the away team in regular time in the second leg increases its probability of winning the penalty shootout by 12.5 percentage points. However, if the home team wins in regular time, it has only a 50% chance to qualify after a shootout. van Hemert and van der Kamp, (2026) consider 409 penalty shootouts between the 2008/09 to 2023/24 seasons, from national competitions (cups, supercups, playoffs for promotion, relegation, qualification) in eight European countries, as well as from international club competitions and tournaments for national teams. Both equalising (the randomly assigned Team A scored the last goal) and conceding (Team A conceded the last goal) dummies are insignificant in their logistic regressions.
@@
Stronger football teams have a significantly higher chance of winning a penalty shootout based on 1067 shootouts from the domestic cup competitions of ten European associations (the top five, as well as Belgium, Portugal, Russia, Turkey, Ukraine), the Brazilian cup, the UEFA Champions League, and the UEFA Europa League (Wunderlich et al.,, 2020). However, the effect of team strength on success remains rather small; the winning probability does not exceed 60% even against an extremely weak team. The novelty of Wunderlich et al., (2020) resides in the measure of strength difference, determined by winning probabilities based on pre-match betting odds, adjusted for home advantage. The recent analysis of all available penalty shootouts until the 2023/24 season reinforces that the favourites identified by pre-match odds are more likely to win (Pipke,, 2025, Table A.5). But the favourite team might not be the stronger team due to home advantage, as remarked by Wunderlich et al., (2020).
@@
van Hemert and van der Kamp, (2026) investigate the effect of team strength quantified by the market value of the starting 11 and the average market value of the players, including the goalkeeper, who participated in the shootout. Based on 409 observations, the winning percentages gradually increase for stronger teams; teams in quartile 4 with the highest relative market value win with a probability of almost 60%. This association is reinforced by logistic regressions: if the market value of a team relative to its opponent increases by 1%, the odds of winning the shootout are higher by 1.3%. On the other hand, none of the covariates—team age, playing at home/away, scoring the last goal, being the first-mover—has a significant impact on winning.
To summarise, all previous studies considering the strength of teams have found at least some bias towards the stronger team in winning penalty shootouts. In other words, penalty shootouts are not equivalent to a pure coin toss; ability is an important success factor even for a mechanical task carried out in a high-pressure environment.
@@ 3 Data
Champions League Europa League Conference League Total Qualification 41 (19) 103 (31) 63 (63) 207 (113) Knockout stage 14 (5) 26 (8) 10 (10) 50 (23) Final 6 (0) 5 (3) 0 (0) 11 (3) Total 61 (24) 134 (42) 73 (73) 268 (139) UEFA Conference League was called UEFA Europa Conference League until the 2023/24 season. Knockout stage stands for the knockout stage without the final. The knockout stage contains two-legged matches, except for the final, which is played on a neutral field. Numbers in parenthesis show the matches played from the 2020/21 season to 2025.
@@
Furthermore, the betting odds published by bookmakers are usually related to the outcome of one particular match, but our database mostly contains the second legs of knockout ties. Therefore, the favourite team identified by betting odds might not be the stronger team because of home advantage. Even though Wunderlich et al., (2020, Section 2.3) offers a method to adjust this bias, which allows the classification of a slight favourite as the slightly weaker team, such a procedure is not necessarily reliable in two-legged matches. The reason is the incentives of the teams: one team could be satisfied by a small loss if it still implies its qualification. For instance, if the away team won the first match by two goals, it does not have powerful incentives to attack in the second match even if it has scored one goal less than the other team. The betting odds of a home/away win certainly reflect the effect of these incentives, which have substantially changed from the 2021/22 season due to the abolition of the away goals rule.
Compared to Elo ratings, betting odds can indeed reflect better the current form of the teams by taking, for example, injuries, into account. However, it is far from clear that these factors are relevant to penalty shootouts, where success is based on individual skills, not on collaboration within a team. The use of Elo ratings, given by an “objective” mathematical formula, may also be more persuasive for decision-makers than betting odds driven by market sentiment. Finally, Football Club Elo Ratings have recently been demonstrated to outperform UEFA club coefficients, the official measure of team strength in these competitions, with respect to predictive accuracy (Csató,, 2024). Based on these arguments, we follow the approach of previous papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025; van Hemert and van der Kamp,, 2026), and do not consider alternative measures of team ability.
@@ 4 Methodology
The connection between the relative strength of teams and the probability of winning the penalty shootout is assessed by logistic regressions, too. The binary dependent variable is the home win/loss. Therefore, matches played on a neutral venue are removed due to the small sample size, similar to Wunderlich et al., (2020). Denote the Elo rating of the home and the away team by $E_{h}$ and $E_{a}$, respectively. The independent variable $\Delta_{ha}$ is either the difference of Elo ratings ($\Delta_{ha}^{(1)}=E_{h}-E_{a}$), or the winning probability of the home team implied by the Elo ratings according to the Elo equation of Football Club Elo Ratings (http://clubelo.com/system):
@@
As a robustness check, two alternative definitions of the comeback team are considered. In the first, $c_{h}=1$ ($c_{h}=-1$) if the home (away) team scores both more goals and the last goal. In the second, $c_{h}$ equals the goal difference from the perspective of the home team, similar to the method of Krumer, (2021).
@@
$P_{h}=\frac{1}{1+e^{-\left(\beta_{0}+\beta_{1}\Delta+\beta_{2}f_{h}+\beta_{3}c_{h}\right)}},$
@@
5 Results
Section 5.1 investigates potential success factors in penalty shootouts by binomial tests, that is, whether teams in a given category win significantly more/less often than 50% compared to teams in the other category. Section 5.2 considers more dimensions simultaneously via logistic regressions. In both cases, sensitivity analyses are also conducted.
@@ 5.1 Partial analysis
Figure 4 focuses on the winning probability of the stronger team based on Football Club Elo Ratings. Five definitions of the favourite are used by gradually restricting the sample according to the minimal difference in the Elo ratings of the two teams. However, even teams that have at least 100 points more than their opponents fail to win 50% of their penalty shootouts. These calculations can be reproduced on the basis of Table A.2 in Appendix A.2, which also reports the associated $p$-values. The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025; van Hemert and van der Kamp,, 2026) demonstrate a significantly higher winning probability for a “sufficiently” stronger team.
As we have seen in Section 3, the number of penalty shootouts has substantially increased from the 2020/21 season. Thus, all binomial tests have also been conducted for the first 20 and the last six seasons, respectively. According to Tables A.1A.2 in the Appendix, the conclusions remain unchanged if we focus on the beginning or the end of the dataset. The only significant value at the level of 5% (but not at the level of 1%) appears for home advantage based on the extended dataset: home teams are less likely to win a penalty shootout than their opponents between 2020 and 2025 if the matches played behind closed doors are taken into account (Table A.1). In these 20 one-legged matches, only seven were won by the home team. This can be explained by choking under pressure (Harb-Wu and Krumer,, 2019), which was probably stronger than usual despite the absence of spectators, as, in contrast to the standard rule, the randomly drawn home team did not play away before. The marginally significant effect disappears when only proper two-legged matches are considered under the restricted definition of home venue (Table A.1). Tables A.3 and A.4 reinforce the insignificant impact of all factors considered by distinguishing the penalty shootouts played in the qualification and knockout phases.